OpenStudy (anonymous):

3^(5x+2)=7^(x-1) How/why is ln used to solve this type of equations??

OpenStudy (anonymous):

Because if you remember, in an equation, in order to mantain the equality, you have to put the inverse of what you got on the other side, a product goes to the other side as a division, a plus changes to a minus and so on. There you have an exponential function so the inverse of the exponential function is the Ln. That's why you use it to solve.

OpenStudy (anonymous):

y = log ax, a^y = x

OpenStudy (amistre64):

it has to do with a change of base procedure

OpenStudy (anonymous):

3^(5x+2)=7^(x-1) Ln(3^5x+2)= ln(7^x-1) (5x+2) ln(3)= (x-1) ln(7) But that as far as I can go.. Idk what to do next? Do I distribute orrrr divide etc.?

OpenStudy (amistre64):

let ln(7)/ln(3) = k therefore 5x+2 = (x-1) * k 5x+2 = kx - k 5x-kx = -2 - k x(5-k) = -2 - k x = -(2+k)/(5-k)

OpenStudy (amistre64):

another way to see that was to go this route: \[3^{5x+2}=7^{x-1}\] \[L_3(7^{x-1})=5x+2\] \[(x-1)L_3(7)=5x+2\] \[L_3(7)=\frac{L(7)}{L(3)}=k\] \[(x-1)k=5x+2\]

OpenStudy (anonymous):

My study guide says the answer is x= -2 log(3) - log(7)/ 5 log(3) - log(7)

OpenStudy (amistre64):

it most likely simplifies to that after alot of heartache :)

OpenStudy (anonymous):

Ugh, too much heartache lol

OpenStudy (amistre64):

\[x = -\frac{(2+k)}{(5-k)}\] \[x = -\frac{(2+\frac73)}{(5-\frac73)}\] \[x = -\frac{(2.3+.7)}{(5-\frac73).3}\] \[x = -\frac{(2.3+.7)\cancel{.3}}{(5.3-.7)\cancel{.3}}\] \[x = \frac{-2.3-.7}{5.3-.7}\]

OpenStudy (anonymous):

How do you differentiate when to use log(x) or ln(x)

OpenStudy (amistre64):

recall that the derivative of log(x) is 1/x say: y = 3x lny = ln(3x) , derive y'/y = 3/(3x) , solve for y' y' = y/x = 3x/x =3

OpenStudy (amistre64):

logx = lnx/ln10 derive = 1/xln(10)